OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (5, 9); Slope: 2

OpenStudy (hoblos):

y-9=2(x-5) y = 2x -1

OpenStudy (anonymous):

y-y1= m(x-x1)

OpenStudy (anonymous):

where m is slope (5,9) = (x1,y1)

OpenStudy (anonymous):

these are my option : y- 9 = 2(x -5) y -5 = 2(x + 9) y + 9 = 2(x -5) y -9 = 2(x + 5)

OpenStudy (hoblos):

so keep it as y- 9 = 2(x -5)

OpenStudy (anonymous):

okay that is what i thought thanks @hoblos